OpenStudy (anonymous):

2x-y+2z=-7 -x+2y-4z=5 x+4y-6z=-1 Solve The System Using The Elimination Method.

OpenStudy (anonymous):

adding eqautions 2 and 3 you elimnate the x and get: 6y-10z=4 (eqn3') mutlt (eqn2) by 2 and add it to (eqn1) to get: 3y-6z=3 (eqn1') now solve the system involving (eqn3') and (eqn1') for y and z by mutliplying (eqn1') by -2, and then adding the two equations together 6y-10z=4 eqn3' -6y+12z=-6 (eqn1'') 2z=-2 z=-1 back sub to get y using (eqn1'): 3y-6(-1)=3 3y=-3 y=-1 now go back to the orginial (eqn3) to get x: x+4(-1)-6(-1)=-1 x-4+6=-1 x=-3 thus the solution set is the ordered triple {(-3, -1, -1)} (assuming no arith. errors!)